OpenStudy (anonymous):

) Suppose line A passes through points (2, -3) and (7, 2) and line B passes through (-3, 7) and (2, 2). Show, algebraically, that lines A and B are perpendicular.

OpenStudy (texaschic101):

first you need to find the slopes by using the slope formula. Then compare the slopes. If they are negative reciprocals of each other, then they are perpendicular

OpenStudy ($oumya):

slope of line A=m1=(2+3)/(7-2)=1 slope of line B=m2=(2-7)/(2+3)=1 we know that, if two straight lines are perpendicular..then.. m1*m2=-1 so here m1*m2 is also -1... so the lines are perpendicular